OpenStudy (anonymous):

Write a revursive and an explicit formula to represent the following arithmetic sequence. 2,-5,-12,-19......

OpenStudy (anonymous):

its 2+0=2 2+(-7)=-5 -5+(-7)=-12 -12+(-7)=-19 means simply add -7